Adult tickets to a play cost $22. Tickets for children cost $15. Tickets for 11 people cost a total of $228.00. Write and solve a system of equations to find how many children and how many adults were in the group
Answer by mananth(16946)

Adult tickets x numbers
child tickets y numbers

22 x + 15 y = 228 .............1
1 x + 1 y = 11 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by -15
22 x 15 y = 228
-15 x -15 y = -165
Add the two equations
7 x = 63
/ 7
x = 9.00 OR 9
plug value of x in (1)
22 x + 15 y = 228
198 + 15 y = 228
15 y = 228-198
15 y = 30
y = 2.00 OR
Adult tickets 9 numbers
child tickets 2 numbers
m.ananth@hotmail.ca